OpenStudy (anonymous):

Frank purchased a boat for $7,525. He made a down payment of $475. He applied for a six-year installment loan with an interest rate of 8.4% in the amount of $7,050. What is the total cost of the boat after six years?

OpenStudy (kropot72):

Interest on the loan for one year = 7050 * 0.084 = $592.20 TotaI interest paid after six years = 592.20 * 6 = $3,553.20 Total cost of the boat after six years =$7,525 + $3,553.20 = $11,078.20

OpenStudy (anonymous):

I just had this question in virtual school, the total amount after 6 years is , $9,474.28